# 3D Mesh Editing using Masked LRMs

### v1

@@ Abstract
We present a novel approach to mesh shape editing, building on recent progress in 3D reconstruction from multi-view images. We formulate shape editing as a conditional reconstruction problem, where the model must reconstruct the input shape with the exception of a specified 3D region, in which the geometry should be generated from the conditional signal. To this end, we train a conditional Large Reconstruction Model (LRM) for masked reconstruction, using multi-view consistent masks rendered from a randomly generated 3D occlusion, and using one clean viewpoint as the conditional signal. During inference, we manually define a 3D region to edit and provide an edited image from a canonical viewpoint to fill in that region. We demonstrate that, in just a single forward pass, our method not only preserves the input geometry in the unmasked region through reconstruction capabilities on par with SoTA, but is also expressive enough to perform a variety of mesh edits from a single image guidance that past works struggle with, while being 10×10$\times$10 × faster than the top-performing competing prior work.
*This work was conducted during an internship at Meta
Figure 1: We present a novel and intuitive approach to perform 3D editing through simple interactions. Our method leverages the representation power of large reconstruction models by learning to inpaint randomly masked multi-view images. At inference time, we provide a single input image edit along with a 3D mask and our model faithfully edits the 3D geometry in just a few seconds.
@@ 1 Introduction
Recent developments in 3D content generation initially followed a similar path to 2D content generation. Operating in 3D voxel space instead of 2D pixel space, models like VAEs [50, 63, 26, 51] and GANs [91, 13] were built, and trained on small-scale datasets [8]. These works often demonstrated limited editing capabilities through simple latent operations on their learned representations. Efforts have been made to extend generative diffusion to 3D as well [49, 92, 75, 43]. There has also been work done in generative autoregressive models for 3D, which tokenize 3D data in a unique way [55, 72, 70, 57]. Furthermore, recent research in neural representation techniques such as Neural Radiance Fields (NeRFs) [53] and Gaussian splatting [38] has introduced an entirely new paradigm for 3D content generation.
@@
Despite significant progress in 3D content generation from scratch, research in editing the shape of existing 3D models is underdeveloped. Image editing methods benefit from a nearly endless source of data from scraping the internet, while 3D assets typically require a higher level of expertise and specialized tools to create and thus are scarce in comparison. The difference in scale is staggering, with the largest image datasets containing billions of samples [65]. In contrast, the largest 3D datasets contains only millions [19].
@@
A common approach to tackling the issue of 3D data scarcity is to exploit existing text-image foundation models. Many recent efforts in 3D editing involve using these huge models to provide guidance to an optimization process by giving them differentiably rendered images of the manipulated geometry as input [52, 54, 25, 5]. While these approaches demonstrated some success, they face several significant challenges. Firstly, the gradients obtained using foundation models as guidance are often extremely noisy, leading to both unstable and unpredictable optimizations [77]. Furthermore, since these methods often use text as input in lieu of visual input, they are hard to control. Finally, these techniques typically directly optimize properties of an explicit 3D mesh, which severely constrains the type of edits that are possible. For example, it is impossible to add a hole to a shape, since such a modification is not topology-preserving.
@@
A different type of approach that many recent works use is to take a two-stage method, placing the brunt of the “creative effort" onto 2D models, using them to generate and edit content. Then, a model trained to lift 2D images into 3D content is used to produce the final output [42, 83]. Thus, by giving the model edited image inputs, a 3D edit is obtained. However, these methods also have their own limitations. While editing a single image is no longer a challenging task, 3D models that lift a single image into geometry often suffer from ambiguous 3D structure in the occluded regions and do not accurately reconstruct what a shape looks like from every viewpoint. Efforts have been made to train multi-view 2D generation models to combat this problem [48, 68]. Unfortunately, as we demonstrate, editing multi-view images in a consistent manner remains a challenging task and can introduce artifacts corrupting the shape construction.
@@
In this paper, we introduce a method that falls into the second aforementioned direction of lifting 2D images into 3D. Our method differs in that instead of using a 3D model to simply reconstruct geometry, our model is inherently trained to “inpaint" regions of multi-view images. Specifically, the inputs to our method are a set of masked renders and a single clean conditional image that is provided to infer the missing information from. Our approach solves the issues present in both approaches to shape editing. In contrast to optimization methods, our model is efficient as it constructs shapes in a single, fast forward pass. Furthermore, the output of our model is highly predictable, as it is trained to reconstruct geometry to a high degree of accuracy. Our approach also solves both the consistency and ambiguity problems existing reconstruction methods have, by relying on only a single conditional image while propagating the conditional signal to the rest of the multi-view inputs.
@@
A key challenge of this approach is designing a training procedure that allows the model to learn how to use the conditional information in a multi-view consistent manner. To accomplish this, we employ a new 3D masking strategy. We mask each input view in a consistent manner by rendering an explicit occluding mesh. Then, by supervising both the occluded and unoccluded regions with multi-view reconstruction targets, our model learns to not only fill in the occluded region, but also to accurately reconstruct the rest of the shape. We demonstrate that this training method allows our model to be used downstream for editing tasks while maintaining strong quantitative performance on reconstruction baselines. By manually defining an editing region analogous to the train-time occlusions, and using a single edited canonical view, users can use our model to generate a shape that is faithful both to the original shape, and the edited content. In summary, our contributions are as follows:
@@
  • We design a novel conditional LRM trained with a new 3D-consistent multi-view masking strategy. Despite not being our primary intention, our architecture matches SoTA reconstruction metrics, while simultaneously learning to use the conditional input to fill in 3D occlusions.
@@
  • We demonstrate that our LRM can be used for 3D shape editing. Our model is much faster than optimization-based editing, and synthesizes arbitrary edits that optimization cannot (e.g. genus changes). Furthermore, our model does not suffer from the multi-view consistency and occlusion ambiguity issues that other LRM-based approaches have.
@@ 2 Related Work
Shape Editing: Editing 3D shapes has been an active area of research for at least four decades. Early works focused on deformation [66, 16], cut and paste [62, 6], Laplacian surface editing [24, 46, 69] or fusion [37]. Recent works have tackled this task from different viewpoints depending on the geometry representation, the losses, and the downstream application. Regarding representation, research has been conducted on implicit surface editing [31, 14], voxels [67], mesh-based deformations [25, 39, 64], NeRF [28, 81, 87, 10, 3, 34, 36] and Gaussian splatting [80, 12, 40]. Another line of work has focused on generative local editing using diffusion models. MagicClay [5] proposed to sculpt 3D meshes using a 2D hybrid representation where part of the geometry is frozen and the remaining is optimized using SDS loss [59, 44]. 3D shape editing has been explored in the context of sketches,  [71, 56, 47], faces [2, 27, 23, 9, 60] or in an interactive manner [20]. Unlike past work, our approach builds upon the recent advancements of LRMs [32, 82] and introduces a novel architecture trained on multi-view consistent masked data that enables 3D mesh editing within a seconds.
@@
Figure 2: Training Pipeline. The images and camera poses (Plücker rays) are patchified and projected into tokens. A random 3D mask is then generated and tokens corresponding to occluded patches are replaced by a learnable mask token. The camera and image tokens are summed and concatenated with learnable triplane tokens to form the transformer input. A clean, posed image serving as the conditional input is also tokenized, forming the cross-attention input. The output triplane tokens are upsampled and decoded into colors and SDF values, which are then transformed into densities for volumetric rendering.
@@
Masked Vision Models: The original Denoising Autoencoder [76] used random noise as a “masking" method on images during training with the objective of learning better semantic representations. More recently, methods using transformers convert images into sequences and predict unknown pixels [4, 11] which culminated in the development of the Vision Transformer (ViT) [21] as the backbone of modern masked image models. Models like the Masked Autoencoder [29] use ViTs to process randomly masked token sequences, where every token represents a distinct, non-overlapping image patch Research in diffusion models which also uses random noise as a “masking" procedure, has exploded in popularity, producing increasingly impressive generated images. By taking random Gaussian noise and constraining it to a specific region, diffusion models can be used for image inpainting [41, 17]. Masked autoencoders have been built for 3D data types such as point clouds [58, 90, 35], meshes [45], and NeRFs [33], with each work developing a different way to “patchify" their respective 3D representations. Point clouds have the most natural representation for next token prediction [70, 57], while efforts have also been made into tokenizing triangle meshes for generation as sequences of vertices and faces [55, 72]. Our paper presents a new approach to combine masking with LRMs for editing.
@@ 3 Method
Our large reconstruction model, shown in Figure 2, reconstructs a 3D shape from input images. Specifically, the model maps a sparse set of renders from various viewpoints into a latent triplane representation. We sample this representation to obtain latent features at different 3D points, which are then decoded into distance and RGB values for volumetric rendering. At training, we predict output renders from arbitrary camera poses. During inference, we use marching cubes to produce the reconstructed geometry. Unlike existing LRMs, our model uses a conditional branch to accept an additional view of the target shape. The inputs are then corrupted by a random masking procedure, forcing the model to learn to “inpaint" the input views using the conditional branch signal.
@@ 3.1 Masked LRM: Architecture
[…]
@@
[…]
@@ 3.3 Training Stages
Figure 3: Addressing Multi-view Artifacts: Editing using a multi-view diffusion model (Zero-123++) is prone to artifacts and produces inaccurate geometry due to ambiguity caused by occlusion. While explicit masking improves results, it still suffers from blurriness artifacts which propagate into the generated shape when such images are used for 3D reconstruction. Our model, which does not use multi-view diffusion, reconstructs the correct geometry.
@@ 3.4 Mesh Shape Editing
Since our LRM is trained with 3D-consistent, multi-view occlusions, using the conditional branch to complete the partial observations, it is straightforward to use it for shape editing. Given a shape $\mathcal{S}$, we manually define an occlusion $\mathcal{O}$ that occludes the region of interest for editing. Then, we edit a representative image within the pixels that are occluded from its camera viewpoint. This may be done a variety of ways – for our results, we use a text-to-image masked diffusion method [41, 17]. The image edit is used as a conditional signal, while the rest of the occluded images are fed to the main transformer body of the LRM. The LRM is trained to inpaint the occluded region using the content from the conditional branch, and as such it propagates the 2D edit into 3D space. This approach to shape editing is much faster than optimization-based methods (see Table 2), requiring only a single forward pass to lift the desired edit into 3D. Our model also produces more realistic shapes since it is trained on a large collection of scanned objects instead of relying on diffusion model guidance and complex regularizations. It is also far more expressive than optimization-based methods as it can generate arbitrary geometry based on the input condition. For example, it can change the geometric genus of a shape (e.g. adding a handle or a hole), which deformation-based optimization methods cannot do as genus changes are not differentiable. Generative methods using LRMs such as InstantMesh [83] rely on a method such as Zero-123++ [68] to generate multi-view images, introducing the risk of view-consistency artifacts. In Figure 3 we provide an example of these artifacts. The original Zero-123++ generates inaccurate views due to the ambiguity of the occluded geometry in the input image. We then restrict Zero-123++ to a masked region in each view, guaranteeing consistency in the rest of the shape (by copying back the original pixels of these regions), but still the model generates severe blurring artifacts. In contrast, our model requires only a single view as conditioning and uses the prior from the dataset to construct the shape in a consistent manner.
@@ 4 Experiments
Training Data. We train our Masked LRM on a 480k subset of the Objaverse dataset [18] containing shapes collected from a wide variety of sources. This subset filters out relatively poor quality 3D models [^1] Sketchfab data were filtered out due their licence. Each shape is normalized to fit within a sphere of fixed radius. Our training data consists of 40 $512\times 512$ images of each shape, rendered from randomly sampled cameras. We also render the corresponding depth and normal maps for these camera poses. Every iteration, the model inputs and reconstruction targets are chosen randomly from these pre-fixed sets of images.
Evaluation. We evaluate the reconstruction quality of our model on the GSO [22] and ABO [15] datasets and compare the state-of-the-art MeshLRM [82]. We use PSNR, SSIM, and LPIPS on the output renders from novel poses as metrics. To remain consistent with the training setting, we randomly generate a rectangular mask to occlude the input views and provide a different clean view as conditioning. Furthermore, we qualitatively demonstrate the main contribution of our model, the ability to propagate 2D edits from a single viewpoint into 3D. We compare our results to prior works for both text and image based 3D generation.
Figure 4: Mesh Editing - Qualitative Comparisons: We compare our approach to other shape editing methods. Given a bird mesh (top left) as well as the various image edits as guidance we demonstrate that our approach is the only one that generates multi-view consistent shape edits that follow the guidance. We omit the colors to clearly visualize the edited geometry.

### v2

@@ Abstract
We present a novel approach to shape editing, building on recent progress in 3D reconstruction from multi-view images. We formulate shape editing as a conditional reconstruction problem, where the model must reconstruct the input shape with the exception of a specified 3D region, in which the geometry should be generated from the conditional signal. To this end, we train a conditional Large Reconstruction Model (LRM) for masked reconstruction, using multi-view consistent masks rendered from a randomly generated 3D occlusion, and using one clean viewpoint as the conditional signal. During inference, we manually define a 3D region to edit and provide an edited image from a canonical viewpoint to fill that region. We demonstrate that, in just a single forward pass, our method not only preserves the input geometry in the unmasked region through reconstruction capabilities on par with SoTA, but is also expressive enough to perform a variety of mesh edits from a single image guidance that past works struggle with, while being $2-10\times$ faster than the top-performing prior work.
@@ 1 Introduction
Recent developments in 3D content generation initially followed a similar path to 2D content generation. Operating in 3D voxel space instead of pixel space, models like VAEs [55, 70, 28, 56] and GANs [98, 14] were built, and trained on small-scale datasets [9]. These works often demonstrated limited editing capabilities through simple latent operations on their learned representations. Efforts have been made to extend generative diffusion to 3D [53, 99, 82, 45]. There has also been work done in generative autoregressive models for 3D, which tokenize 3D data in a unique way [60, 79, 77, 62]. Furthermore, neural representation techniques such as NeRFs [58] and Gaussian splatting [40] have introduced an entirely new paradigm for 3D content generation.
@@
Despite significant progress in 3D content generation from scratch, research in editing the shape of existing 3D models is underdeveloped. Image editing methods benefit from a nearly endless source of data from scraping the internet, while 3D assets typically require a higher level of expertise and specialized tools to create and thus are scarce in comparison. The difference in scale is staggering, with the largest image datasets containing billions of samples [72] while the largest 3D datasets contain only millions [20].
[…]
Figure 1: Mesh Editing using MaskedLRMs: The inputs comprise front/back views of the source mesh $(1^{st}$ column) and a frontal image used as the conditional view. The $2^{nd}$ column shows the masked area rendered from the front (inset) and a 2D edit. The last column shows our generated mesh from the front/back views.
@@
A common approach to tackling the issue of 3D data scarcity is to exploit existing text-image foundation models. Recent efforts in 3D editing involve using these huge models to provide guidance to an optimization process by giving them differentiably rendered images of the manipulated geometry as input [57, 59, 27, 6]. While these approaches demonstrated some success, they face several major challenges. Firstly, the gradients obtained using foundation models as guidance are often extremely noisy, leading to unstable and unpredictable optimizations [84]. Furthermore, since these methods often use text as input in lieu of visual input, they are hard to control. Finally, these techniques typically directly optimize properties of an explicit 3D mesh, which severely constrains the type of possible edits. For example, it is impossible to add a hole to a shape, since such a modification is not topology-preserving.
@@
Recent works follow a different path and utilize a two-stage approach, placing the brunt of the “creative effort” onto 2D models, using them to generate and edit content. Then, a pipeline that lifts 2D images into 3D content produces the final output [44, 90]. Thus, by giving the model edited image inputs, a 3D edit is obtained. However, these methods rely on diffusion models that produce multi-view images [50, 75, 54, 52] which then are passed to a 3D reconstruction model [89, 34]. While editing a single image is no longer a challenging task, this multi-view generation procedure often suffers from ambiguous 3D structure in the occluded regions and does not accurately reconstruct what a shape looks like from every viewpoint. Efforts have been made to adapt multi-view diffusion models specifically for text-driven editing instead of the single-view-to-multi-view task [5, 24]. As we qualitatively demonstrate, editing multi-view images in a realistic manner remains a challenging task.
@@
Our proposed approach falls into the second direction: lifting 2D images to 3D. Instead of using a 3D model to simply reconstruct geometry, our model is inherently trained to “inpaint” regions of multi-view images. The inpainting task is performed directly in 3D, instead of in multi-view image space. Specifically, the inputs to our method are a set of masked renders and a single clean conditional image that is provided to infer the missing information from. Our approach solves the issues present in both approaches to shape editing. In contrast to optimization methods, our model is efficient as it constructs shapes in a single, fast forward pass. Furthermore, the output of our model is highly predictable, as it is trained to reconstruct and inpaint geometry to a high degree of accuracy. This predictability gives a high degree of control to our method via the conditioning image. Our approach addresses the multi-view consistency and ambiguity problems of reconstruction methods by relying on a single conditional image while propagating the conditional signal to the rest of the multi-view inputs.
@@
A key challenge is designing a training procedure that allows the model to learn how to use the conditional information in a multi-view consistent manner. To accomplish this, we introduce a new 3D masking strategy. We mask each input view in a consistent manner by rendering an explicit occluding mesh. Then, by supervising both the occluded and unoccluded regions with multi-view reconstruction targets, our model learns to not only fill in the occluded region, but also to accurately reconstruct the rest of the shape. Unlike previous works such as NeRFiller [88] which used fixed masks at a per-scene basis, training with randomly generated masks allows our model to generalize to arbitrary shapes and test-time masks. We demonstrate that this training method allows our model to be used downstream for editing tasks while maintaining strong quantitative performance on reconstruction baselines. By manually defining an editing region analogous to the train-time occlusions, and using a single edited canonical view, users can use our model to generate a shape that is faithful both to the original shape, and the edited content. In summary, our contributions are as follows:
  • We design a novel conditional LRM trained with a new 3D-consistent multi-view masking strategy that enables our LRM to generalize to arbitrary masks during inference.
@@
  • Despite not being our primary intention, our architecture matches SoTA reconstruction metrics, while concurrently learns to use the conditional input to fill 3D occlusions.
@@
  • We show that our LRM can be used for 3D shape editing while being $2-10\times$ faster than optimization- and LRM-based edit methods. It synthesizes edits that optimization cannot (e.g. genus changes) and does not suffer from the multi-view consistency and occlusion ambiguity issues that approaches trained without masking suffer from.
@@ 2 Related Work
Figure 2: Training Pipeline. The images and camera poses are patchified and projected into tokens. A random 3D mask is generated and tokens corresponding to occluded patches are replaced by a learnable mask token. Camera and image tokens are summed and concatenated with learnable triplane tokens to form the transformer input. A clean conditional image is tokenized, forming the cross-attention input. The output triplane tokens are upsampled and decoded into colors and SDF values, which are transformed into densities for volumetric rendering.
Shape Editing: Editing 3D shapes has been an active area of research for at least four decades. Early works focused on deformation [73, 17], cut and paste [68, 7], Laplacian surface editing [26, 48, 76] or fusion [39]. Recent works have tackled this task from different viewpoints depending on the geometry representation, the losses, and the downstream application. Regarding representation, research has been conducted on implicit surface editing [33, 15], voxels [74], mesh-based deformations [27, 41, 71], NeRF [30, 88, 94, 11, 3, 36, 38] and Gaussian splatting [87, 13, 42, 67]. Another line of work focused on generative local editing using diffusion models. MagicClay [6] sculpted 3D meshes using a 2D hybrid representation where part of the geometry is frozen and the remaining is optimized using SDS loss [64, 46]. 3D shape editing has been explored in the context of sketches,  [78, 61, 49], faces [2, 29, 25, 10, 65] or in an interactive manner [21]. Recent approaches build upon progress in LRMs [34, 89], performing multi-view editing using diffusion models and then using LRMs to reconstruct an edited shape [5, 66, 24]. In contrast, our work introduces a novel architecture trained on multi-view consistent masked data that bypasses the need for inconsistent diffusion editing and enables 3D mesh editing within seconds.
@@
Masked Vision Models: The original Denoising Autoencoder [83] used random noise as a “masking” method on images during training with the objective of learning better semantic representations. More recently, methods using transformers convert images into sequences and predict unknown pixels [4, 12] which culminated in the development of the Vision Transformer (ViT) [22] as the backbone of modern masked image models. Models like the Masked Autoencoder [31] use ViTs to process randomly masked tokens, where every token represents a distinct, non-overlapping image patch Research in diffusion models which also uses random noise as a “masking” procedure, has exploded in popularity, producing increasingly impressive generated images. By taking random Gaussian noise and constraining it to a specific region, diffusion models can be used for image inpainting [43, 18]. Masked autoencoders have been built for 3D data types such as point clouds [63, 97, 37], meshes [47], and NeRFs [35], with each work developing a different way to “patchify” their respective 3D representations. Point clouds have the most natural representation for next token prediction [77, 62], while efforts have also been made into tokenizing triangle meshes for generation as sequences of vertices and faces [60, 79]. Our paper presents a new approach to combine masking with LRMs for editing.
@@ 3 Method
Our large reconstruction model, shown in Figure 2, reconstructs a 3D shape from input images. Specifically, the model maps a sparse set of renders from various viewpoints into a latent triplane representation. We sample this representation to obtain latent features at different 3D points, which are then decoded into distance and RGB values for volumetric rendering. At training, we predict output renders from arbitrary camera poses. During inference, we use marching cubes to produce the reconstructed geometry. Unlike existing LRMs, our model uses a conditional branch to accept an additional view of the target shape. The inputs are then corrupted by a random masking procedure, forcing the model to learn to “inpaint” the input views using the conditional branch signal.
@@ 3.1 Masked LRM: Architecture
$T_{\text{Image}}^{i}=\textbf{PatchEmbed}(\mathbf{I}_{i}),T_{\text{Plücker}}^{i}=\textbf{PatchEmbed}(\mathbf{C}_{i}),$ (1)
where PatchEmbed is the operation of splitting images into non-overlapping patches and applying a linear layer to the channels. We reserve $T_{\text{Image}}^{1}$ and $T_{\text{Plücker}}^{1}$ for the clean conditional signal. Now, we sample a random rectangular mesh $\mathcal{O}$ and render it from the same camera poses as $\mathcal{S}$. Comparing the depth maps of $\mathcal{S}$ and $\mathcal{O}$, we produce modified tokens $\tilde{T}_{\text{Image}}^{i}$ where the token for any patch that contains an occluded pixel is replaced by a learnable mask token. Then, the input tokens to the model are constructed as:
$\mathbf{x}=\underset{i=2}{\overset{n}{\Big{\|}}}(\tilde{T}_{\text{Image}}^{i}+T_{\text{Plücker}}^{i}),\;\;\;\mathbf{z}=T_{\text{Image}}^{1}+T_{\text{Plücker}}^{1},$ (2)
@@
$\sigma$ $=\textbf{Density}(\textbf{SDF})$
$\hat{\mathbf{N}}$ $=\textbf{NormGrad}(\textbf{SDF})$
@@ 3.3 Training Stages
Figure 3: Multi-view Diffusion vs MaskedLRM: Multi-view diffusion models must infer occluded geometry from the input which leads to artifacts in the multi-view images (incorrect/distorted views). MaskedLRM which does not use multi-view diffusion, receives both the edited image and multi-view information, allowing it to bypass this problem and reconstruct the correct geometry.
@@ 3.4 Mesh Shape Editing
Since our LRM is trained with 3D-consistent, multi-view occlusions, using the conditional branch to complete the partial observations, it is straightforward to use it for shape editing. Given a shape $\mathcal{S}$, we manually define an occlusion $\mathcal{O}$ that occludes the region of interest for editing. Then, we edit a representative image within the pixels that are occluded from its camera viewpoint. This may be done a variety of ways – for our results, we use a text-to-image masked diffusion method [43, 18]. The image edit is used as a conditional signal, while the rest of the occluded images are fed to the main transformer body of the LRM. The LRM is trained to inpaint the occluded region using the content from the conditional branch, and as such it propagates the 2D edit into 3D space. This approach to shape editing is much faster than optimization-based methods (see Table 2), requiring only a single forward pass to lift the desired edit into 3D. Our model also produces more realistic shapes since it is trained on a large collection of scanned objects instead of relying on diffusion model guidance and complex regularizations. It is also more expressive than optimization-based methods as it can generate arbitrary geometry based on the input condition. For example, it can change the geometric genus of a shape (adding a handle or a hole as in Figure 4), which deformation-based optimization methods cannot do as genus changes are not differentiable. Generative methods using LRMs such as InstantMesh [90] rely on methods such as Zero-123++ [75] to generate multi-view images, introducing view-consistency artifacts. In Figure 3 we show examples of such artifacts generated by recent models. Zero-123++ hallucinates additional holes in the vase, and generates a distorted and incorrect bird anatomy. SyncDreamer [52] generates unrealistically distorted views, such as a completely flattened vase, poor bird anatomy, and a warped chair. Wonder3D [54] is better, but it cannot capture the correct bird anatomy and chair structure. In contrast, our model requires only a single view as conditioning and uses the prior from the dataset to construct the shape in a consistent manner. Some recent concurrent work tackles editing directly in the multi-view image space. While this also handles ambiguity, we show in Figure 6 that our method produces more realistic edits.
@@ 4 Experiments
Training Data. We train our Masked LRM on a the Objaverse dataset [19] containing shapes collected from a wide variety of sources. Each shape is normalized to fit within a sphere of fixed radius. Our training data consists of 40 $512\times 512$ images of each shape, rendered from randomly sampled cameras. We also render the corresponding depth and normal maps for these camera poses. Every iteration, the model inputs and reconstruction targets are chosen randomly from these pre-fixed sets of images.
Evaluation. We evaluate the reconstruction quality of our model on the GSO [23] and ABO [16] datasets and compare the state-of-the-art MeshLRM [89]. Since MeshLRM cannot be easily repurposed for our editing task, we also compare the reconstruction quality with InstantMesh. We use PSNR, SSIM, and LPIPS on the output renders from novel poses as metrics. To remain consistent with the training setting, we randomly generate a rectangular mask to occlude the input views and provide a different clean view as conditioning for our method. Finally, we qualitatively demonstrate the main contribution of our model, the ability to propagate 2D edits from a single viewpoint into 3D. We compare our results to prior works for text and image based 3D generation.